OpenStudy (haltiamali):

Whoever gets this question will get a fan and a medal!!! In 1999, there were 9860 Great Danes registered with the American Kennel Club. The number of registered Labrador Retrievers was 6997, more than fifteen times the number registered Great Danes. How many registered Labrador Retrievers were there? Show me how you did it, please!

OpenStudy (anonymous):

i think there is a misplaced comma here

OpenStudy (anonymous):

x=6997+15*9860

OpenStudy (anonymous):

SOlve for x

OpenStudy (anonymous):

Can u?

OpenStudy (haltiamali):

Um..

OpenStudy (anonymous):

x=6997+15*9860 x=6997+147900 x=154897

OpenStudy (haltiamali):

Thanks.

OpenStudy (anonymous):

Welcome